A motorist drives 5.5km from his home to work in 35 minutes but makes the return trip in only 25 minutes. What are the average speeds and average veloceties for:





(1) each half of the round trip





(2) total trip​

1
Expert's answer
2021-10-31T18:08:47-0400

1)


"v_1=5.5\frac{60}{35}=9.4\frac{km}{h}\\
v_2=5.5\frac{60}{25}=13.2\frac{km}{h}\\"

2)


"v=2\cdot5.5\frac{60}{35+25}=11\frac{km}{h}"

Average velocity is zero.